OpenStudy (anonymous):

the function of f(x) is -x^2 - x - 6 what is f(2) A. -8 B. 0 C. x - 10 D. x^2 - 4

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hi.

OpenStudy (misty1212):

\[ f(x) = -x^2 - x - 6\\ f(2) = -2^2 - 2 - 6\]

OpenStudy (anonymous):

Ohhh.

OpenStudy (anonymous):

So, 0?